OpenStudy (anonymous):

What is the slppe of 3x+4y=12

OpenStudy (anonymous):

slope?

OpenStudy (anonymous):

To find the slope of the line, put in slope-intercept form or y=, so: subtract -3x 4y=-3x+12 divide by 4 y=-4/3x+3 The slope is -4/3.

OpenStudy (anonymous):

zeinas can you help me again?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I posted a question

OpenStudy (anonymous):

Tag me in it, so I can find it easier

OpenStudy (shamil98):

\[\huge ax + by = c\] in standard form, you can easily determine the slope by taking those values and using the formula: \[\huge m = -\frac{ a }{ b }\] where m = slope